Community Coffee Company wants a new flavor of Cajun Coffee. How many pounds of cofee worth $7 a pound should be added to 14 pounds of coffee worth $4 a pound to get a mixture worth $5 a pound?
And there is this chart i have to fill in

Answer by checkley77(12844)

7x+14*4=5(x+14)
7x+56=5x+70
7x-5x=70-56
2x=14
x=14/2
x=7 pounds of 7$ coffee is used.
Proof:
7*7+14*4=5(7+14)
49+56=5*21
105=105